# Undefinability of Approximation of 2-to-2 Games

### v1

@@ Abstract
Recent work by Atserias and Dawar [6] and Tucker-Foltz [28] has established undefinability results in fixed-point logic with counting ($\mathrm{FPC}$) corresponding to many classical complexity results from the hardness of approximation. In this line of work, $\mathrm{NP}$-hardness results are turned into unconditional $\mathrm{FPC}$ undefinability results. We extend this work by showing the $\mathrm{FPC}$ undefinability of any constant factor approximation of weighted $2$-to-$2$ games, based on the $\mathrm{NP}$-hardness results of Khot, Minzer and Safra. Our result shows that the completely satisfiable $2$-to-$2$ games are not $\mathrm{FPC}$-separable from those that are not $\epsilon$-satisfiable, for arbitrarily small $\epsilon$. The perfect completeness of our inseparability is an improvement on the complexity result, as the $\mathrm{NP}$-hardness of such a separation is still only conjectured. This perfect completeness enables us to show the $\mathrm{FPC}$ undefinability of other problems whose NP-hardness is conjectured. In particular, we are able to show that no $\mathrm{FPC}$ formula can separate the $3$-colourable graphs from those that are not $t$-colourable, for any constant $t$.
@@ 1 Introduction
Perhaps the most important open question in the field of the hardness of approximation is the unique games conjecture of Khot. This states that for any $\epsilon,\delta>0$, there is a set of labels $\Sigma$ such that it is $\mathrm{NP}$-hard to separate the $(1-\epsilon)$-satisfiable instances of $\Sigma$-unique games (the precise definitions follow below) from those that are not even $\delta$-satisfiable. The strongest result obtained so far in this direction shows that there is a $\Sigma$ for which it is $\mathrm{NP}$-hard to separate the $(\frac{1}{2}-\epsilon)$-satisfiable instances from the $\delta$-unsatisfiable ones. This result is a consequence of the $2$-to-$2$ theorem due to Khot, Minzer and Safra [21, 11, 22].
@@
In the present paper, we consider the $\mathrm{FPC}$ definability of $2$-to-$2$ games. The hardness of approximating the optimum value of such games was established through a series of results by Khot, Minzer and Safra [21, 27, 11]. At the core of their proof is a reduction from the problem MAX 3XOR of maximizing the number of satisfied clauses in a 3XOR instance. We show that the reductions used can be formulated, with some modification, as first-order definable reductions. As a consequence, we obtain the result that the completely satisfiable instances of $2$-to-$2$ games cannot be separated by an $\mathrm{FPC}$ formula from those that are no more than $\delta$-satisfiable, for arbitrarily small $\delta$. This $(1,\delta)$ separation is stronger (in terms of approximation ratios) than the known $(1-\epsilon,\delta)$ $\mathrm{NP}$-hardness result due to the fact that the $\mathrm{FPC}$ undefinability of approximating MAX 3XOR was proved with perfect completeness in [6]. A corollary of our result is the $\mathrm{FPC}$ undefinability of a $(\frac{1}{2},\delta)$ separation for unique games. This improves, again in terms of the approximation ratios, the gap obtained by Tucker-Foltz, though it should be noted that the latter gap is for a more restriced class of games.
@@
The result on graph colouring should be compared with a recent result of Atserias and Dalmau [5] which shows that the promise graph colouring problem cannot be solved by a local consistency algorithm. In particular, this implies that for any constant $t$ the $3$-colourable graphs cannot be separated from those that are not $t$-colourable by a class (whose complement is) definable in Datalog. Since Datalog programs can be translated into sentences of $\mathrm{FPC}$, our Theorem 5.3 can be seen as strengthening their result. It is worth examining this relationship more closely. It is known, from results of [4] and [8], that every class of bounded counting width (and therefore, in particular, any $\mathrm{FPC}$ definable class) that is the complement of a fixed-template constraint satisfaction problem (CSP) is already definable in Datalog. Hence, we can conclude from the result of Atserias and Dalmau that no $\mathrm{FPC}$ definable CSP separates the $3$-colourable graphs from the non-$t$-colourable ones. However, since it is conceivable that a separating class for these two CSP is $\mathrm{FPC}$ definable but not itself a CSP, our result is still a strengthening. But we can say still more. It can be deduced from the proof in [5] that the $3$-colourable graphs and the non-$t$-colourable ones are not separable by any class definable in an existential positive infinitary logic ($\exists^{+,\omega}$). Moreover, it is a consequence of a very recent proof due to Rossman [26] that every class of bounded counting width that is preserved under homomorphisms is definable in $\exists^{+,\omega}$. Thus, we can conclude from these results that no homomorphism-closed class of bounded counting width separates the $3$-colourable graphs from the non-$t$-colourbale ones. Since Theorem 5.3 easily applies to all classes of bounded counting width and not just the $\mathrm{FPC}$-definable ones; and it is conceivable that a separating class is not necessarily closed under homomorphisms, our result subsumes even this strengthened version of that of Atserias and Dalmau.
@@ 2.1 Hardness of Approximation in Optimization
We are interested in $\mathrm{NP}$-hard optimization problems. A standard example is the problem MAX 3SAT, where the aim is to find, given a formula in 3CNF, an assignment of values to its variables that maximizes the number of clauses satisfied. Formally, consider a function problem $M$, which associates with every possible input instance $I$ a value $M(I)$. In our example, MAX 3SAT maps a formula $\phi$ to the maximum number $m$ of clauses of $\phi$ that can be simultaneously satisfied. While, in practice, we might be interested in finding an assignment that achieves this maximum, for the purpose of proving hardness, it suffices to show that it is hard to compute the number $m$. When finding $M(I)$ is hard, we may wish to approximate it, and we say that an algorithm computes a $C$-approximation (for a real number $C>1$) of $M$ if it produces a number $M^{\prime}(I)$ with the guarantee that $M^{\prime}(I)\leq M(I)\leq C\cdot M^{\prime}(I)$.
@@
Formally, let $A$ and $B$ be two sets (i.e. decision problems) with $A\cap B=\emptyset$. We say that $A$ and $B$ are $\mathrm{NP}$-hard to separate, if every set $C$ with $A\subseteq C\subseteq\overline{B}$ is $\mathrm{NP}$-hard, where $\overline{B}$ denotes the complement of $B$. For a function problem $M$, and a constant $c\in[0,1]$, denote by $c$-$M$ the set $\{I\mid M(I)\geq c\}$. Then, for constants $c$ and $s$ with $0\leq s<c\leq 1$, we say that the gap problem $\text{Gap}M(c,s)$ is $\mathrm{NP}$-hard if it is $\mathrm{NP}$-hard to separate the sets $c$-$M$ and $\overline{s\text{-}M}$. This implies, in particular, that unless $\mathrm{P}=\mathrm{NP}$, there is no polynomial-time algorithm giving a $\frac{c}{s}$-approximation of $M$. The value $c$ in $\text{Gap}M(c,s)$ is called the completeness parameter and $s$ the soundness parameter
[…]
@@
A common way of deriving further hardness of approximation results is via gap-reductions: given function problems $A$ and $B$, a polynomial-time computable function $f$ taking instances of $A$ to instances of $B$ is a reduction from $\mathrm{Gap}A(c,s)$ to $\mathrm{Gap}B(c^{\prime},s^{\prime})$ if for all instances $I$ of $A$
@@ 2.2 Label Cover Games
A game $(G,\Sigma,\Phi)$ is simple if it is both edge consistent and edge distinct.
Note that in a simple game, there is at most one edge between any pair of vertices. Most of the games we consider with multiple edges are edge consistent. We mention edge distinct and simple games mainly to constrast with other results in the literature. For an edge consistent game $(G,\Sigma,\Phi)$, and an edge $e$ between vertices $u$ and $v$, we sometimes write $\Phi(u,v)$ instead of $\Phi(e)$. No ambiguity arises since by definition of edge consistency, the value of $\Phi$ is the same on all edges between $u$ and $v$.
In this paper, we are particularly interested in $2$-to-$2$ games and $1$-to-$1$ games, the latter also being known as Unique Games. We write $\mathrm{UG}_{q}$ for the function problem of determining the value of an instance of unique games with an alphabet of size $q$. We can then state Khot’s unique games conjecture.
@@
For any $\delta,\epsilon$ with $10<\delta<1-\epsilon<$ 1, there exists a positive integer $q$ so that $\mathrm{Gap}\mathrm{UG}_{q}(1-\epsilon,\delta)$ is NP-Hard.
@@
The best known hardness result for unique games, towards proving Conjecture 2.2 is that $\mathrm{GapUG}_{q}(\frac{1}{2}-\epsilon,\delta)$ is $\mathrm{NP}$-Hard for arbitrarily small $\delta$ and $\epsilon$. This is obtained as a consequence of the hardness of $2$-to-$2$ games established by Khot, Minzer and Safra, which we return to in Section 3.
@@
[…]
@@
We are also concerned with weighted $2$-to-$2$ and $1$-to-$1$ games, attaching a weight to each constraint.
@@
[…]
@@ 2.3 Undefinability of Approximation
For relational structures $\mathbb{A}$ and $\mathbb{B}$ in the same vocabulary, and a positive integer $k$, $\mathbb{A}\equiv^{k}\mathbb{B}$ denotes that the two structures cannot be distinguished by any sentence of first-order logic with counting using no more than $k$ distinct variables. For a class $\mathcal{C}$ of structures, the counting width of $\mathcal{C}$ is the function $\nu:\mathbb{N}\rightarrow\mathbb{N}$ such that for any $n$, $\nu(n)$ is the least $k$ such that $\mathcal{C}$, restricted to structures with at most $n$ elements is a union of $\equiv^{k}$-equivalence classes. Any class that is definable by a sentence of $\mathrm{FPC}$ has counting width bounded by a constant. Almost all results showing that a class is not definable in $\mathrm{FPC}$ proceed by showing that it, in fact, does not have bounded counting width.
@@ 3 The Reduction
The proof of Theorem 2.3 was completed in 2018 and remains to this day the most significant advance towards establishing the Unique Games Conjecture since the latter was formulated by Khot in [19]. The proof proceeds by a reduction from $\mathrm{Gap}\textsc{3XOR}(1-\epsilon,\frac{1}{2}+\delta)$ and was presented in a series of papers [21, 27, 11]. The main difficulty lies in proving the combinatorial conditions that the soundness analysis relies on. The full reduction and proof of correctness can be found in [24, Chapter 3].
@@
For every $\delta$ with $10<\delta<$ 1, there exists $q\in\mathbb{N}^{+}$ for which $\mathrm{Gap}\textsc{2to2}_{q}(1,\delta)$ is not $\mathrm{FPC}$ definable.
@@ 3.1 Regular 3XOR
[…]
@@ 3.2 Reducing to Transitive Games
In the first step of the reduction, we reduce regular 3XOR instances to label cover games with a mixture of $2$-to-$2$ and $1$-to-$1$ constraints, with an additional transitivity requirement. We formally define these below.
@@
An edge consistent game $(G,\Sigma,\Phi)$ is a transitive 2-to-2 game if $\Phi:E\rightarrow\mathcal{P}(\Sigma^{2})$ assigns to each edge $e$ either a $2$-to-$2$ or a $1$-to-$1$ relation and whenever $\Phi(u,v)$ is $1$-to-$1$, then for any edge $(v,w)$, $\Phi(u,w)$ is the composition of $\Phi(u,v)$ and $\Phi(v,w)$.
Note that the condition on composition only applies when $\Phi(u,v)$ is $1$-to-$1$, but $\Phi(v,w)$ may be $1$-to-$1$ or $2$-to-$2$, and this determines whether $\Phi(u,w)$ is $1$-to-$1$ or $2$-to-$2$.
@@
For $U=(e_{1},\ldots,e_{k})\in\mathcal{U}$, let $X_{U}$ denote the set of variables occuring in equations in $U$ and for $i\in\{1,\ldots,k\}$ let $v_{i}\in\mathbb{F}_{2}^{X}$ denote the vector which has $1$s in the three coordinates corresponding to the variables occurring in $e_{i}$ and 00s everywhere else. We define the space of side-conditions corresponding to $U$ to be $H_{U}=\mathrm{Span}(v_{1},\dots,v_{k})$. Note that, by construction, the vectors are $v_{1},\ldots,v_{n}$ are linearly independent and thus form a basis for $H_{U}$. We say that a linear function $f:\mathbb{F}_{2}^{X}\rightarrow\mathbb{F}_{2}$ satisfies the equations in $U$ if $f(v_{i})=b_{i}$ for all $i$, where $b_{i}$ is the right-hand side of $e_{i}$.
@@
We can now define the reduction $\Theta$ that takes the instance $I$ to a $2$-to-$2$ transitive game $\Theta(I)$. The reduction depends on the choice of parameters $k$ and $l$. We omit the details on how to select the right parameters.
@@
Given a pair of vertices $u=(U,L)$ and $v=(U^{\prime},L^{\prime})$, the constraint $\Phi(u,v)$ is a $1$-to-$1$ relation if
@@
It is the case for any pair, that $\mathrm{dim}(L+H_{U}+H_{U^{\prime}})=\mathrm{dim}(L^{\prime}+H_{U}+H_{U^{\prime}})$ [21, Lemma 4.3]. Let us call this dimension $D$. By [21, Lemma 4.4], any linear function $f:L+H_{U}\rightarrow\mathbb{F}_{2}$ satisfying the equations of $U$ has a unique extension to $(L+H_{U}+H_{U^{\prime}})$ that also satisfies the equations of $U^{\prime}$. Then, it is easily seen that if $\mathrm{dim}(L+L^{\prime}+H_{U}+H_{U^{\prime}})=D$, then $f$ has exactly one label of $(U^{\prime},L^{\prime})$ that it is consistent with, and if $\mathrm{dim}(L+L^{\prime}+H_{U}+H_{U^{\prime}})=D+1$, there are exactly two such functions, thanks to the “free dimension”. Hence, the constraints are $1$-to-$1$ or $2$-to-$2$ as required. The transitivity property of these constraints is established in [21, Appendix A].

### v2

# arXiv:2504.03523v2[cs.LO] 01 Mar 2026

@@ Abstract
Recent work by Atserias and Dawar [7] and Tucker-Foltz [29] has established undefinability results in fixed-point logic with counting ($\mathrm{FPC}$) corresponding to many classical complexity results from the hardness of approximation. In this line of work, $\mathrm{NP}$-hardness results are turned into unconditional $\mathrm{FPC}$ undefinability results. We extend this work by showing the $\mathrm{FPC}$ undefinability of any constant factor approximation of weighted 22-to-22 games, based on the $\mathrm{NP}$-hardness results of Khot, Minzer and Safra. Our result shows that the completely satisfiable 22-to-22 games are not $\mathrm{FPC}$-separable from those that are not $\epsilon$-satisfiable, for arbitrarily small $\epsilon$. The perfect completeness of our inseparability is an improvement on the complexity result, as the $\mathrm{NP}$-hardness of such a separation is still only conjectured. This perfect completeness enables us to show the $\mathrm{FPC}$ undefinability of other problems whose NP-hardness is conjectured. In particular, we are able to show that no $\mathrm{FPC}$ formula can separate the $3$-colourable graphs from those that are not $t$-colourable, for any constant $t$.
@@ 1 Introduction
Perhaps the most important open question in the field of the hardness of approximation is the unique games conjecture of Khot. This states that for any $\epsilon,\delta>0$, there is a set of labels $\Sigma$ such that it is $\mathrm{NP}$-hard to separate the $(1-\epsilon)$-satisfiable instances of $\Sigma$-unique games (the precise definitions follow below) from those that are not even $\delta$-satisfiable. The strongest result obtained so far in this direction shows that there is a $\Sigma$ for which it is $\mathrm{NP}$-hard to separate the $(\frac{1}{2}-\epsilon)$-satisfiable instances from the $\delta$-unsatisfiable ones. This result is a consequence of the 22-to-22 theorem due to Khot, Minzer and Safra [24, 13, 28].
@@
In the present paper, we consider the $\mathrm{FPC}$ definability of 22-to-22 games. The hardness of approximating the optimum value of such games was established through a series of results by Khot, Minzer and Safra [24, 28, 13]. At the core of their proof is a reduction from the problem MAX 3XOR of maximizing the number of satisfied clauses in a 3XOR instance. We show that the reductions used can be formulated, with some modification, as first-order definable reductions. As a consequence, we obtain the result that the completely satisfiable instances of 22-to-22 games cannot be separated by an $\mathrm{FPC}$ formula from those that are no more than $\delta$-satisfiable, for arbitrarily small $\delta$. This $(1,\delta)$ separation is stronger (in terms of approximation ratios) than the known $(1-\epsilon,\delta)$ $\mathrm{NP}$-hardness result due to the fact that the $\mathrm{FPC}$ undefinability of approximating MAX 3XOR was proved with perfect completeness in [7] (that is to say, with completeness parameter $1$). A corollary of our result is the $\mathrm{FPC}$ undefinability of a $(\frac{1}{2},\delta)$ separation for unique games. This improves, again in terms of the approximation ratios, the gap obtained by Tucker-Foltz, though it should be noted that the latter gap is for a more restriced class of games.
@@
The result on graph colouring should be compared with a recent result of Atserias and Dalmau [6] which shows that the promise graph colouring problem cannot be solved by a local consistency algorithm. In particular, this implies that for any constant $t$ the $3$-colourable graphs cannot be separated from those that are not $t$-colourable by a class (whose complement is) definable in Datalog. Since Datalog programs can be translated into sentences of $\mathrm{FPC}$, our Theorem 5.3 can be seen as strengthening their result. It is worth examining this relationship more closely. It is known, from results of [5] and [9], that every class of bounded counting width (and therefore, in particular, any $\mathrm{FPC}$ definable class) that is the complement of a fixed-template constraint satisfaction problem (CSP) is already definable in Datalog. Hence, we can conclude from the result of Atserias and Dalmau that no $\mathrm{FPC}$ definable CSP separates the $3$-colourable graphs from the non-$t$-colourable ones. However, since it is conceivable that a separating class for these two CSP is $\mathrm{FPC}$ definable but not itself a CSP, our result is still a strengthening. Indeed, in the context of PCSP, the results of [6] imply that $\mathrm{FPC}$ can define more than Datalog can. In particular they give an example of a PCSP that is not solvable by local consistency methods but is solvable by a Sherali-Adams lift of the linear programming formulation. Since we know [12] that such lifts of linear programming formulations of CSP are obtained as first-order interpretations and that the solvability of linear programs is expressible in $\mathrm{FPC}$ [2], it follows that $\mathrm{FPC}$ is strictly more powerful than Datalog for PCSP separations. Thus, our result on graph colouring is indeed a strengthening.
@@ 2.1 Hardness of Approximation in Optimization
We are interested in $\mathrm{NP}$-hard optimization problems. A standard example is the problem
MAX 3SAT, where the aim is to find, given a formula in 3CNF, an assignment of values to its variables that maximizes the number of clauses satisfied. Formally, consider a function problem $M$, which associates with every possible input instance $I$ a value $M(I)$. In our example, MAX 3SAT maps a formula $\phi$ to the maximum number $m$ of clauses of $\phi$ that can be simultaneously satisfied. While, in practice, we might be interested in finding an assignment that achieves this maximum, for the purpose of proving hardness, it suffices to show that it is hard to compute the number $m$. When finding $M(I)$ is hard, we may wish to approximate it, and we say that an algorithm computes a $C$-approximation (for a real number $C>1$) of $M$ if it produces a number $M^{\prime}(I)$ with the guarantee that $M^{\prime}(I)\leq M(I)\leq C\cdot M^{\prime}(I)$.
@@
Formally, let $A$ and $B$ be two sets (i.e. decision problems) with $A\cap B=\emptyset$. We say that $A$ and $B$ are $\mathrm{NP}$-hard to separate [^1] Note that this definition is, despite appearances, symmmetric with respect to $A$ and $B$ if we understand $\mathrm{NP}$-hardness to mean hardness under Turing reuctions., if every set $C$ with $A\subseteq C\subseteq\overline{B}$ is $\mathrm{NP}$-hard, where $\overline{B}$ denotes the complement of $B$. For a function problem $M$, and a constant $c\in[0,1]$, denote by $c$-$M$ the set $\{I\mid M(I)\geq c\}$. Then, for constants $c$ and $s$ with $0\leq s<c\leq 1$, we say that the gap problem $\text{Gap}M(c,s)$ is $\mathrm{NP}$-hard if it is $\mathrm{NP}$-hard to separate the sets $c$-$M$ and $\overline{s\text{-}M}$. This implies, in particular, that unless $\mathrm{P}=\mathrm{NP}$, there is no polynomial-time algorithm giving a $\frac{c}{s}$-approximation of $M$. The value $c$ in $\text{Gap}M(c,s)$ is called the completeness parameter and $s$ the soundness parameter
The first hardness of approximation results come from the PCP theorem [3, 17, 4]: one of its direct consequences is the $\mathrm{NP}$-hardness of $\mathrm{Gap}\textsc{3SAT}(1,\eta)$ for some constant $\eta$ strictly less than $1$. Håstad [20] obtained an optimum inapproximability result for MAX 3SAT. Namely, he showed that $\mathrm{Gap}\textsc{3SAT}(1,\frac{7}{8}+\epsilon)$ is $\mathrm{NP}$-hard for arbitrarily small $\epsilon$. This is optimal since every 3CNF is $\frac{7}{8}$-satisfiable. Similarly, he also showed that $\mathrm{Gap}\textsc{3XOR}(1-\epsilon,\frac{1}{2}+\epsilon)$ is $\mathrm{NP}$-hard for arbitrarily small $\epsilon$. Again, this is optimal. Here, 3XOR is the problem where we are given a Boolean formula as a conjunction of clauses, each of which is the XOR of three literals and we aim to maximize the number of satisfied clauses. Note that the completeness parameter must be strictly less than $1$, since the problem of determining whether such a formula is satisfiable or not is polynomial-time decidable. Thus 11-3XOR can be separated in polynomial time from $\overline{(1-\epsilon)\text{-}\textsc{3XOR}}$ for any $\epsilon>0$.
@@
A common way of deriving further hardness of approximation results is via gap-reductions: given function problems $A$ and $B$, a polynomial-time computable function $f$ taking instances of $A$ to instances of $B$ is a reduction from $\mathrm{Gap}A(c,s)$ to $\mathrm{Gap}B(c^{\prime},s^{\prime})$ if for all instances $I$ of AA
@@ 2.2 Label Cover Games
A game $(G,\Sigma,\Phi)$ is simple if it is both edge consistent and edge distinct. In other words, there is at most one edge between any pair of vertices.
Most of the games we consider with multiple edges are edge consistent. We mention edge distinct and simple games mainly to constrast with other results in the literature. For an edge consistent game $(G,\Sigma,\Phi)$, and an edge $e$ between vertices $u$ and $v$, we sometimes write $\Phi(u,v)$ instead of $\Phi(e)$. No ambiguity arises since by definition of edge consistency, the value of $\Phi$ is the same on all edges between $u$ and $v$.
In this paper, we are particularly interested in 22-to-22 games and 11-to-11 games, the latter also being known as Unique Games. We write $\mathrm{UG}_{q}$ for the function problem of determining the value of an instance of unique games with an alphabet of size $q$. We can then state Khot’s unique games conjecture.
@@
For any $\delta,\epsilon$ with $0<\delta<1-\epsilon<1$, there exists a positive integer $q$ so that $\mathrm{Gap}\mathrm{UG}_{q}(1-\epsilon,\delta)$ is NP-Hard.
@@
The best known hardness result for unique games, towards proving Conjecture 2.2 is that $\mathrm{GapUG}_{q}(\frac{1}{2}-\epsilon,\delta)$ is $\mathrm{NP}$-Hard for arbitrarily small $\delta$ and $\epsilon$. This is obtained as a consequence of the hardness of 22-to-22 games established by Khot, Minzer and Safra, which we return to in Section 3.
@@
For any $\delta,\epsilon$ with $0<\delta<1-\epsilon<1$, there exists a positive integer $q$ so that $\mathrm{Gap2to2}_{q}(1-\epsilon,\delta)$ is NP-Hard.
@@
We are also concerned with weighted 22-to-22 and 11-to-11 games, attaching a weight to each constraint.
@@
We write $\mathcal{WG}_{2:2;q}$ to denote the class of weighted 22-to-22 games with $q$ labels and $\textsc{Weight2to2}_{q}$ to denote the function taking such a game to its value. Similarly, we write $\textsc{UG}_{q}$ and $\textsc{WeightUG}_{q}$ for the functions giving the values of unique games and weighted unique games with $q$ labels respectively.
@@ 2.3 Undefinability of Approximation
For relational structures $\mathbb{A}$ and $\mathbb{B}$ in the same vocabulary, and a positive integer $k$, $\mathbb{A}\equiv_{C^{k}}\mathbb{B}$ denotes that the two structures cannot be distinguished by any sentence of first-order logic with counting using no more than $k$ distinct variables. For a class $\mathcal{C}$ of structures, the counting width of $\mathcal{C}$ is the function $\nu:\mathbb{N}\rightarrow\mathbb{N}$ such that for any $n$, $\nu(n)$ is the least $k$ such that $\mathcal{C}$, restricted to structures with at most $n$ elements is a union of $\equiv_{C^{k}}$-equivalence classes. Any class that is definable by a sentence of $\mathrm{FPC}$ has counting width bounded by a constant. Almost all results showing that a class is not definable in $\mathrm{FPC}$ proceed by showing that it, in fact, does not have bounded counting width. Note that the Weisfeiler-Leman dimension of a graph $G$ is defined to be the smallest $k$ such that for any graph $H$ that is not isomorphic to $G$, we have $G\not\equiv_{C^{k+1}}H$. This is a graph parameter associating a number with a graph in contrast to counting width (as defined in [12]) which associates a function with a class $\mathcal{C}$.
@@ 3 The Reduction
The proof of Theorem 2.3 was completed in 2018 and remains to this day the most significant advance towards establishing the Unique Games Conjecture since the latter was formulated by Khot in [21]. The proof proceeds by a reduction from $\mathrm{Gap}\textsc{3XOR}(1-\epsilon,\frac{1}{2}+\delta)$ and was presented in a series of papers [24, 28, 13]. The main difficulty lies in proving the combinatorial conditions that the soundness analysis relies on. The full reduction and proof of correctness can be found in [26, Chapter 3].
@@
For every $\delta$ with $0<\delta<1$, there exists $q\in\mathbb{N}^{+}$ for which $\mathrm{Gap}\textsc{2to2}_{q}(1,\delta)$ is not $\mathrm{FPC}$ definable.
@@ 3.1 Regular 3XOR
An instance of 3XOR can be seen as a system of linear equations over the field $\mathbb{F}_{2}$ with exactly three variables appearing in each equation. We say that such an instance is $d$-regular if every variable appears in at most $d$ equations and no two equations share more than one variable. It is known that for some constant $\frac{1}{2}<\eta<1$, the $\mathrm{NP}$-hardness of $\mathrm{Gap}\textsc{3XOR}(1-\epsilon,\eta)$ holds even when restricted to $d$-regular instances for some fixed value of $d$ (indeed, taking $d=5$ suffices, see [26, Theorem 3.3.1]). In Section 4.3 we show that this is also true of the undefinability in $\mathrm{FPC}$ of $\mathrm{Gap}\textsc{3XOR}(1,\eta)$ From now on, we restrict attention to $d$-regular instances for a suitable fixed value of $d$, and we call the resulting function problem $\mathrm{Gap}\textsc{Regular3XOR}$.
@@ 3.2 Reducing to Transitive Games
In the first step of the reduction, we reduce regular 3XOR instances to label cover games with a mixture of 22-to-22 and 11-to-11 constraints, with an additional transitivity requirement. We formally define these below.
@@
An edge consistent game $(G,\Sigma,\Phi)$ is a transitive 2-to-2 game if $\Phi:E\rightarrow\mathcal{P}(\Sigma^{2})$ assigns to each edge $e$ either a 22-to-22 or a 11-to-11 relation and whenever $\Phi(u,v)$ is 11-to-11, then for any edge $(v,w)$, $\Phi(u,w)$ is the composition of $\Phi(u,v)$ and $\Phi(v,w)$.
Note that the condition on composition only applies when $\Phi(u,v)$ is 11-to-11, but $\Phi(v,w)$ may be 11-to-11 or 22-to-22, and this determines whether $\Phi(u,w)$ is 11-to-11 or 22-to-22.
@@
For $U=(e_{1},\ldots,e_{k})\in\mathcal{U}$, let $X_{U}$ denote the set of variables occuring in equations in $U$ and for $i\in\{1,\ldots,k\}$ let $v_{i}\in\mathbb{F}_{2}^{X}$ denote the vector which has $1$s in the three coordinates corresponding to the variables occurring in $e_{i}$ and 0s everywhere else. We define the space of side-conditions corresponding to $U$ to be $H_{U}=\mathrm{Span}(v_{1},\dots,v_{k})$. Note that, by construction, the vectors $v_{1},\ldots,v_{k}$ are linearly independent and thus form a basis for $H_{U}$. We say that a linear function $f:\mathbb{F}_{2}^{X}\rightarrow\mathbb{F}_{2}$ satisfies the equations in $U$ if $f(v_{i})=b_{i}$ for all $i$, where $b_{i}$ is the right-hand side of $e_{i}$.
@@
We can now define the reduction $\Theta$ that takes the instance $I$ to a 22-to-22 transitive game $\Theta(I)$. We use parameters $k$ and $l$ throughout the construction, which only depend on $\epsilon$ and $\delta$ (for the purpose of the reduction, they are constants). Their exact dependence on $\epsilon$ and $\delta$ can be found in [26].
@@
Given a pair of vertices $u=(U,L)$ and $v=(U^{\prime},L^{\prime})$, the constraint $\Phi(u,v)$ is a 11-to-11 relation if
@@
It is the case for any pair, that $\mathrm{dim}(L+H_{U}+H_{U^{\prime}})=\mathrm{dim}(L^{\prime}+H_{U}+H_{U^{\prime}})$ [24, Lemma 4.3]. Let us call this dimension $D$. By [24, Lemma 4.4], any linear function $f:L+H_{U}\rightarrow\mathbb{F}_{2}$ satisfying the equations of $U$ has a unique extension to $(L+H_{U}+H_{U^{\prime}})$ that also satisfies the equations of $U^{\prime}$. Then, it is easily seen that if $\mathrm{dim}(L+L^{\prime}+H_{U}+H_{U^{\prime}})=D$, then $f$ has exactly one label of $(U^{\prime},L^{\prime})$ that it is consistent with, and if $\mathrm{dim}(L+L^{\prime}+H_{U}+H_{U^{\prime}})=D+1$, there are exactly two such functions, thanks to the “free dimension”. Hence, the constraints are 11-to-11 or 22-to-22 as required. The transitivity property of these constraints is established in [23, Appendix A]. A more detailed explanation on this can also be found in  [23, Appendix A].